# Provably Convergent Plug-and-play Proximal Block Coordinate Descent Method for Hyperspectral Anomaly Detection

### v1

@@ Abstract
Key words:
Low-rank representation, proximal block coordinate descent, hyperspectral anomaly detection, plug-and-play.
@@ 1 Introduction
In this paper, we develop a novel approach for hyperspectral anomaly detection that integrates a representation-based technique with a deep learning denoiser. We express the background of HSIs using a subspace representation, employ a deep denoiser in a plug-and-play (PnP) fashion to eliminate the noise from the representation coefficients, and identify anomalies using a group sparsity measure. Our main contributions are summarized as follows:
  • We propose an optimization model to detect anomalous objects in HSIs that may be contaminated with noise. The background of HSIs is approximated by a tensor mode-$3$ product of a learnable orthogonal basis as the subspace and a tensor formed by eigenimages as the representation coefficients. To remove noise, we utilize a deep implicit prior with a weakly convex potential. For detecting sparse anomalous objects, we introduce a generalized group sparsity measure $\|\cdot\|_{2,\psi}$, where $\psi$ is a sparsity-promoting function. The resulting model is a nonconvex nonsmooth minimization problem with an orthogonal constraint.
  • We enhance the existing relaxed proximal denoiser, which has a gradient step form. To denoise the eigenimages that may not fall in the pretrained range, we propose a shifted proximal denoiser and formulate its potential using the potential of the existing relaxed proximal denoiser.
  • We introduce a PnP version of the proximal block coordinate descent algorithm, called the PnP-PBCD method, for solving the proposed model. The subproblems for updating the orthogonal basis and the sparse anomalous objects have either closed-form solutions or are easy to compute. We prove that any accumulation point of the sequence generated by the proposed algorithm is a stationary point.
@@ 2.1 Notations
For a third order tensor $\mathcal{X}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, we let $x_{i_{1}i_{2}i_{3}}$ denote its $(i_{1},i_{2},i_{3})$-th entry, let $x_{i_{1}i_{2}:}$ denote its $(i_{1},i_{2})$-th mode-3 fiber and let $X_{::i_{3}}$ denote its $i_{3}$-th frontal slice. The mode-$k$ unfolding of a third order tensor $\mathcal{X}$ is denoted as $\mathcal{X}_{(k)}={\rm unfold}_{(k)}(\mathcal{X})$, which is the process to linearize all indexes except index $k$. The dimensions of $\mathcal{X}_{(k)}$ are $I_{k}\times\prod^{3}_{j=1,j\neq k}I_{j}$. An element $x_{i_{1}i_{2}i_{3}}$ of $\mathcal{X}$ corresponds to the position of $(i_{k},j)$ in matrix $\mathcal{X}_{(k)}$, where $j=1+\sum^{3}_{l=1,l\neq k}(i_{l}-1)\prod^{l-1}_{m=1,m\neq k}I_{m}$. And the inverse process of the mode-$k$ unfolding of a tensor $\mathcal{X}$ is denoted by $\mathcal{X}={\rm fold}_{(k)}(\mathcal{X}_{(k)})$. In particular, the mode-$3$ product of a tensor $\mathcal{Z}\in\mathbb{R}^{n_{1}\times n_{2}\times r}$ and a matrix $Y\in\mathbb{R}^{n_{3}\times r}$, denoted by $\mathcal{Z}\times_{3}Y$, is a tensor $\mathcal{X}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ with entries
@@ 2.2 Formulation for anomaly detection in noisy HSIs
According to the high spectral correlation of HSIs, a clean HSI $\mathcal{L}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ can be expressed in a low-rank tensor representation as follows
$\mathcal{L}=\mathcal{Z}\times_{3}E,$ (1)
where $E\in\mathbb{R}^{n_{3}\times r}$ represents a basis of the spectral subspace and the tensor $\mathcal{Z}\in\mathbb{R}^{n_{1}\times n_{2}\times r}$ denotes the representation coefficient of $\mathcal{L}$ with respect to $E$. In particular, we choose $E$ as an orthogonal basis, that is, $E^{\top}E={I}_{r}$ with ${I}_{r}$ denoting the identity matrix of size $r$. And each band of $\mathcal{Z}$ is called as an eigenimage. Then a noisy HSI $\mathcal{O}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ can be formulated mathematically as
@@
where $\|\cdot\|_{2,\psi}$ represents a generalized group sparsity measure for detecting anomalous objects, $\Phi_{\Sigma}(\cdot)$ represents a proximal denoiser prior for removing Gaussian noise, and $E$ is a learnable orthogonal basis. The resulting model is a nonconvex nonsmooth minimization problem with an orthogonal constraint. More details of this model will be provided in the following subsections.
@@ 2.3 Generalized group sparsity measure
    Minimax concave penalty (MCP) [33]: for $\theta>\gamma$,
@@ 2.4 Proximal denoiser prior in a PnP framework
$g_{\sigma}(X)=\frac{1}{2}\|X-\mathcal{N}_{\sigma}(X)\|^{2},$ (6)
with $\mathcal{N}_{\sigma}(X)$ being a $\mathcal{C}^{2}$ neural network, specifically DRUNet [34], pre-trained for denoising grayscale or color images. And the denoiser is carefully trained to ensure that $g_{\sigma}$ approximately has an $L_{g_{\sigma}}$-Lipschitz gradient with $L_{g_{\sigma}}<1$. The overall denoiser $\mathcal{D}_{\sigma}$ is called a proximal denoiser, because it behaves like a proximal operator as shown in Proposition 2.1. More discussions on this proximal denoiser can also be found in [11, 29, 25].
@@
$\phi_{\sigma}(X)=\begin{cases}g_{\sigma}(\mathcal{D}^{-1}_{\sigma}(X))-\frac{1}{2}\|\mathcal{D}^{-1}_{\sigma}(X)-X\|_{F}^{2},\quad&\mbox{if }X\in\operatorname{Im}(\mathcal{D}_{\sigma})\\ \phi_{\sigma}(X)=+\infty,\quad&\mbox{otherwise}.\end{cases}$ (7)
@@
$\widetilde{\mathcal{D}}_{\sigma}^{\gamma}(Z)=\frac{1}{a}\left[\mathcal{D}_{\sigma}^{\gamma}(aZ+b)-b\right],$
@@
where $\lambda>0$ and $\Sigma:=\operatorname{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{n})$ with $\sigma_{n}$ denoting the noise level of the $n$-th eigenimage ${Z}_{::n}$.
@@ 3 Plug-and-play Proximal Block Coordinate Descent Method
Let $F(\mathcal{Z},E,\mathcal{S})$ denote the objective function of the proposed model (2) and
$H(\mathcal{Z},E,\mathcal{S})=\frac{\delta}{2}\|\mathcal{Z}\times_{3}E+\mathcal{S}-\mathcal{O}\|_{F}^{2}.$ (9)
@@ 3.1 The update of $\mathcal{S}$
$\mathcal{S}^{k+1}\in\operatorname{prox}_{\tilde{\gamma}\|\cdot\|_{2,\psi}}\left[\mathcal{S}^{k}-\tilde{\alpha}_{\mathcal{S}}\left(\mathcal{S}^{k}+\mathcal{Z}^{k}\times_{3}E^{k}-\mathcal{O}\right)\right],$ (13)
where $\tilde{\gamma}=\frac{\gamma}{\delta+\alpha_{\mathcal{S}}}$ and $\tilde{\alpha}_{\mathcal{S}}=\frac{\delta}{\delta+\alpha_{\mathcal{S}}}$. Since $\|\cdot\|_{2,\psi}$ is separable, the $(i,j)$-th mode-$3$ fiber of $\mathcal{S}^{k+1}$ can be computed via
$s^{k+1}_{ij:}\in\operatorname{prox}_{\tilde{\gamma}\psi\circ\|\cdot\|_{2}}(\hat{s}^{k}_{ij:}),$
@@
$\operatorname{prox}_{\tilde{\gamma}\psi\circ\|\cdot\|_{2}}(s)=\begin{cases}\operatorname{prox}_{\tilde{\gamma}\psi}(\|s\|_{2})\frac{s}{\|s\|_{2}},\quad&\|s\|_{2}\neq 0\\ 0,\quad&\|s\|_{2}=0.\end{cases}$
Depending on the choice of $\psi$, the proximal operator of $\operatorname{prox}_{\tilde{\gamma}\psi\circ\|\cdot\|_{2}}$ are computed differently.
@@ 3.2 The update of $E$
$\mathcal{L}=\mathcal{Z}\times_{3}E\quad\mbox{ if and only if }\quad\mathcal{L}_{(3)}=E\mathcal{Z}_{(3)}.$ (16)
Then the function $H$ can be rewritten as $H(\mathcal{Z},E,\mathcal{S})=\frac{\delta}{2}\|E\mathcal{Z}_{(3)}+(\mathcal{S}-\mathcal{O})_{(3)}\|_{F}^{2}$, which have the same form as (14a). It follows from Lemma 3.1 that minimizing (14a) is equivalent to minimizing more simple forms as (14b) and (14c). In particular, we consider $\widetilde{H}(\mathcal{Z},E,\mathcal{S})=-\delta\langle E,(\mathcal{O}-\mathcal{S})_{(3)}(\mathcal{Z}_{(3)})^{\top}\rangle$ of the form as (14b). Then the update of $E$ can be computed as follows
@@
where $\tilde{\alpha}_{E}=\frac{\delta}{\alpha_{E}}$ and $\operatorname{Proj}_{\mathbb{S}_{n_{3},r}}$ denotes the projection onto the Stiefel manifold $\mathbb{S}_{n_{3},r}$. And it follows from Lemma 3.1 in [19] that the projection $\operatorname{Proj}_{\mathbb{S}_{n_{3},r}}$ has a closed form and $E^{k+1}$ can be computed as follows
@@ 3.3 The update of $\mathcal{Z}$
Then the subproblem for updating $\mathcal{Z}$ can be reformulated as
@@
${Z}^{k+1}_{::n}=\widetilde{\mathcal{D}}_{\sigma_{n}}^{\gamma}(\widehat{Z}^{k}_{::n}),$ (20)
@@ 4 Convergence Analysis of the PnP-PBCD Method
In this section, we conduct a convergence analysis on the proposed PnP-PBCD method. We first provide some preliminaries on the subdifferentials for nonconvex nonsmooth functions and the Riemannian gradient of a smooth function on the Stiefel manifold. Then we define the first-order optimality condition of problem (2) and prove that any accumulation point of the sequence generated by the PnP-PBCD method given in Algorithm 1 is a stationary point of problem (2).
4.1 Preliminaries
@@
4.2 First-order optimality condition
We define the first-order optimality condition of the orthogonal constrained optimization problem as in (2). The point $(\bar{\mathcal{Z}},\bar{E},\bar{\mathcal{S}})$ is a first-order stationary point of problem (2) if $0\in\partial F(\bar{\mathcal{Z}},\bar{E},\bar{\mathcal{S}})$, that is,
@@ 4.3 Subsequence convergence
    $\Phi_{\Sigma}$ is coercive and $\rho_{1}$-weakly convex.
@@
Let $f:\mathbb{R}^{d}\to(-\infty,+\infty]$ be a $\rho$-weakly convex function, and let $h:\mathbb{R}^{d}\to\mathbb{R}$ be a differentiable and $\mu$-strongly convex function. If we have
@@
where $\mu-\rho+\alpha>0$, then
$f(x_{0})+h(x_{0})\geq f(\hat{x})+h(\hat{x})+\frac{\mu-\rho+\alpha}{2}\|x_{0}-\hat{x}\|^{2}.$ (24)
@@
Since $f(x)+\frac{\rho}{2}\|x-x_{0}\|^{2}$ and $h(x)-\frac{\mu}{2}\|x-x_{0}\|^{2}$ are convex, we have $f(x)+h(x)+\frac{\alpha}{2}\|x-x_{0}\|^{2}$ is $(\mu-\rho+\alpha)$-strongly convex. Then (23) is well defined and $\hat{x}$ is uniquely obtained.
Next, for the $\rho$-weakly convex function $f$, it follows from Lemma 2.1 in [4] that
@@
for $\forall u\in\partial f(\hat{x})$. And for the differentiable and $\mu$-strongly convex function $h$, it follows from [21] that
@@
By the update of $\hat{x}$ in (23), we have $0\in\partial f(\hat{x})+\nabla h(\hat{x})+\alpha(\hat{x}-x_{0})$. That is, $-\nabla h(\hat{x})+\alpha(x_{0}-\hat{x})\in\partial f(\hat{x})$. Substituting $u=-\nabla h(\hat{x})+\alpha(x_{0}-\hat{x})$ into the above inequality, we obtain (24). ∎
@@
Let $X\in\mathbb{R}^{n\times r}$, $A\in\mathbb{R}^{r\times m}$ and $B\in\mathbb{R}^{n\times m}$. And let $H_{1}(X)$, $H_{2}(X)$, and $H_{3}(X)$ denote the objective functions of (14a), (14b), and (14c), respectively. Then we have
@@
$\nabla H_{1}$ $=(XA-B)A^{\top}=XAA^{\top}+\nabla H_{2}$
$\nabla H_{2}$ $=-BA^{\top}$
$\nabla H_{3}$ $=X-BA^{\top}=X+\nabla H_{2}.$
Then by projecting the gradients above onto the tangent space of Stiefel manifold at $X$, we can compute the Riemannian gradients of $H_{1}$, $H_{2}$, and $H_{3}$. Since $\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{m,n}}$ is linear, we have
$\operatorname{grad}H_{1}(X)$ $=\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{m,n}}(XAA^{\top})+\operatorname{grad}H_{2}(X)$
@@
$\operatorname{grad}H_{3}(X)$ $=\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{m,n}}(X)+\operatorname{grad}H_{3}(X).$
It is easy to verify that
$\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{m,n}}(XAA^{\top})=XAA^{\top}-\frac{1}{2}X(X^{\top}XAA^{\top}+(XAA^{\top})^{\top}X)=0$
and $\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{m,n}}(X)=X-\frac{1}{2}X(X^{\top}X+X^{\top}X)=0$. Therefore, we obtain (25). ∎
@@
Assume that assumptions (A1) and (A2) are satisfied. Let $\{(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\}$ be the sequence generated by Algorithm 1 with $\alpha_{\mathcal{Z}}>\rho_{1}-\delta$ and $\alpha_{\mathcal{S}}>\rho_{2}-\delta$. Then the following statements hold:
@@
$F(\mathcal{Z}^{k-1},E^{k-1},\mathcal{S}^{k-1})-F(\mathcal{Z}^{k-1},E^{k-1},\mathcal{S}^{k})\geq\frac{\delta-\rho_{1}+\alpha_{\mathcal{S}}}{2}\|\mathcal{S}^{k}-\mathcal{S}^{k-1}\|^{2}_{F}$
@@
$F(\mathcal{Z}^{k-1},E^{k},\mathcal{S}^{k})-F(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\geq\frac{\delta-\rho_{2}+\alpha_{\mathcal{Z}}}{2}\|\mathcal{Z}^{k}-\mathcal{Z}^{k-1}\|^{2}_{F}.$

### v3

@@ Abstract
Keywords: Low-rank representation, proximal block coordinate descent, hyperspectral anomaly detection, plug-and-play.
2000 Mathematics Subject Classification: L15A18, 68U10, 46N10.
@@ 1 Introduction
In this paper, we develop a novel approach for hyperspectral anomaly detection that utilizes a representation-based technique for expressing the background, a deep learning denoiser for reducing noise contamination and a group sparsity measure for identifying anomalies. Our main contributions are summarized as follows:
  • We represent the background of HSIs in terms of a tensor mode-$3$ product of a learnable orthogonal basis as the subspace and a tensor formed by eigenimages as the representation coefficients.
  • We employ a deep learning denoiser in a plug-and-play (PnP) fashion to eliminate the noise from the eigenimages. We enhance the existing relaxed proximal denoiser to its shifted version to denoise the eigenimages that may not fall within the pretrained range. The proposed denoiser can also be viewed as a proximal operator associated with a weakly convex function.
  • We introduce a generalized group sparsity measure, $\|\cdot\|_{2,\psi}$, to detect sparse anomalous objects. The function $\psi$ is a sparsity-promoting function and can be chosen as a weakly convex function.
  • We propose a PnP version of the proximal block coordinate descent algorithm, called the PnP-PBCD method, for solving the proposed nonconvex nonsmooth minimization problem with an orthogonal constraint. The subproblems have either closed-form solutions or are easy to compute. We prove that any accumulation point of the sequence generated by the proposed algorithm is a stationary point.
@@ 2.1 Notations
For a third order tensor $\mathcal{X}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, we let $x_{i_{1}i_{2}i_{3}}$ denote its $(i_{1},i_{2},i_{3})$-th entry, let $x_{i_{1}i_{2}:}$ denote its $(i_{1},i_{2})$-th mode-3 fiber and let $X_{::i_{3}}$ denote its $i_{3}$-th frontal slice. The mode-$k$ unfolding of a third order tensor $\mathcal{X}$ is denoted as $\mathcal{X}_{(k)}={\rm unfold}_{(k)}(\mathcal{X})$, which is the process to linearize all indexes except index $k$. The dimensions of $\mathcal{X}_{(k)}$ are $n_{k}\times\prod^{3}_{j=1,j\neq k}n_{j}$. An element $x_{i_{1}i_{2}i_{3}}$ of $\mathcal{X}$ corresponds to the position of $(i_{k},j)$ in matrix $\mathcal{X}_{(k)}$, where $j=1+\sum^{3}_{l=1,l\neq k}(i_{l}-1)\prod^{l-1}_{m=1,m\neq k}n_{m}$. The inverse process of the mode-$k$ unfolding of a tensor $\mathcal{X}$ is denoted by $\mathcal{X}={\rm fold}_{(k)}(\mathcal{X}_{(k)})$. In particular, the mode-$3$ product of a tensor $\mathcal{Z}\in\mathbb{R}^{n_{1}\times n_{2}\times r}$ and a matrix $Y\in\mathbb{R}^{n_{3}\times r}$, denoted by $\mathcal{Z}\times_{3}Y$, is a tensor $\mathcal{X}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ with entries
@@
The expression $\mathcal{X}=\mathcal{Z}\times_{3}Y$ can also be written in terms of unfolding of tensors, i.e., $\mathcal{X}_{(3)}=Y\mathcal{Z}_{(3)}$.
@@ 2.2 Formulation for anomaly detection in noisy HSIs
According to the high spectral correlation of HSIs, a clean HSI can be expressed in a low-rank tensor representation. Specifically, for $\mathcal{L}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, where $n_{1}$ and $n_{2}$ are the spatial dimensions, and $n_{3}$ is the spectral dimension, $\mathcal{L}$ can be represented as follows
$\mathcal{L}=\mathcal{Z}\times_{3}E,$
where $E\in\mathbb{R}^{n_{3}\times r}$ represents a basis of the spectral subspace, and the tensor $\mathcal{Z}\in\mathbb{R}^{n_{1}\times n_{2}\times r}$ denotes the representation coefficient of $\mathcal{L}$ with respect to $E$. In particular, we choose $E$ as an orthogonal basis, that is, $E^{\top}E={I}_{r}$ with ${I}_{r}$ denoting the identity matrix of size $r$. Additionally, each band of $\mathcal{Z}$, denoted as $Z_{::n}$, is called as an eigenimage, where $n=1,2,\dots,r$. Then a noisy HSI $\mathcal{O}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ can be formulated mathematically as
@@
where $\|\cdot\|_{2,\psi}$ represents a generalized group sparsity measure for detecting anomalous objects, $\Phi_{\Sigma}(\cdot)$ represents a proximal denoiser prior for removing Gaussian noise, $E$ is a learnable orthogonal basis, and $\delta,\tau>0$ are parameters. The resulting model is a nonconvex nonsmooth minimization problem with an orthogonal constraint. More details of this model will be provided in the following subsections.
@@ 2.3 Generalized group sparsity measure
    Minimax concave penalty (MCP) [33]: for $\theta>\lambda$,
@@ 2.4 Proximal denoiser prior in a PnP framework
$g_{\sigma}(X)=\frac{1}{2}\|X-\mathcal{N}_{\sigma}(X)\|^{2},$
with $\mathcal{N}_{\sigma}(X)$ being a $\mathcal{C}^{2}$ neural network, specifically DRUNet [34], pre-trained for denoising grayscale or color images. Moreover, the denoiser is carefully trained to ensure that $g_{\sigma}$ approximately has an $L_{g_{\sigma}}$-Lipschitz gradient with $L_{g_{\sigma}}<1$. The overall denoiser $\mathcal{D}_{\sigma}$ is called a proximal denoiser, because it behaves like a proximal operator as shown in Proposition 2.1. More discussions on this proximal denoiser can also be found in [11, 25, 29].
@@
$\phi_{\sigma}(X)=\begin{cases}g_{\sigma}(\mathcal{D}^{-1}_{\sigma}(X))-\frac{1}{2}\|\mathcal{D}^{-1}_{\sigma}(X)-X\|^{2},\quad&\mbox{if }X\in\operatorname{Im}(\mathcal{D}_{\sigma})\\ +\infty,\quad&\mbox{otherwise}.\end{cases}$
@@
$\widetilde{\mathcal{D}}_{\sigma}^{\gamma}(Z)=\frac{1}{a}\left[\mathcal{D}_{\sigma}^{\gamma}(aZ+b)-b\right],$ (5)
@@
where $\lambda>0$ and $\Sigma:=\operatorname{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{r})$ with $\sigma_{n}$ denoting the noise level of the $n$-th eigenimage ${Z}_{::n}$.
@@ 3 Plug-and-play Proximal Block Coordinate Descent Method
Let $F(\mathcal{Z},E,\mathcal{S})$ denote the objective function of the proposed model (1) as follows
$F(\mathcal{Z},E,\mathcal{S})=H(\mathcal{Z},E,\mathcal{S})+\tau\|\mathcal{S}\|_{2,\psi}+\Phi_{\Sigma}(\mathcal{Z}),$ (7)
where
$H(\mathcal{Z},E,\mathcal{S})=\frac{\delta}{2}\|\mathcal{Z}\times_{3}E+\mathcal{S}-\mathcal{O}\|_{F}^{2}.$
@@ 3.1 The update of $\mathcal{S}$
$\mathcal{S}^{k+1}\in\operatorname{prox}_{\tilde{\tau}\|\cdot\|_{2,\psi}}\left[\mathcal{S}^{k}-\tilde{\alpha}_{\mathcal{S}}\left(\mathcal{S}^{k}+\mathcal{Z}^{k}\times_{3}E^{k}-\mathcal{O}\right)\right],$ (11)
where $\tilde{\tau}=\frac{\tau}{\delta+\alpha_{\mathcal{S}}}$ and $\tilde{\alpha}_{\mathcal{S}}=\frac{\delta}{\delta+\alpha_{\mathcal{S}}}$. Since $\|\cdot\|_{2,\psi}$ is separable, the $(i,j)$-th mode-$3$ fiber of $\mathcal{S}^{k+1}$ can be computed via
$s^{k+1}_{ij:}\in\operatorname{prox}_{\tilde{\tau}\psi\circ\|\cdot\|_{2}}(\hat{s}^{k}_{ij:}),$
@@
$\operatorname{prox}_{\tilde{\tau}\psi\circ\|\cdot\|_{2}}(s)=\begin{cases}\operatorname{prox}_{\tilde{\tau}\psi}(\|s\|_{2})\frac{s}{\|s\|_{2}},\quad&\|s\|_{2}\neq 0\\ 0,\quad&\|s\|_{2}=0.\end{cases}$
Depending on the choice of $\psi$, the proximal operator $\operatorname{prox}_{\tilde{\tau}\psi\circ\|\cdot\|_{2}}$ is computed differently.
@@ 3.2 The update of $E$
$\mathcal{L}=\mathcal{Z}\times_{3}E\quad\mbox{ if and only if }\quad\mathcal{L}_{(3)}=E\mathcal{Z}_{(3)}.$
Then the function $H$ can be rewritten as $H(\mathcal{Z},E,\mathcal{S})=\frac{\delta}{2}\|E\mathcal{Z}_{(3)}+(\mathcal{S}-\mathcal{O})_{(3)}\|_{F}^{2}$, which has the same form as (12a). It follows from Lemma 3.1 that minimizing (12a) is equivalent to minimizing more simple forms as (12b) and (12c). In particular, we consider $\widetilde{H}(\mathcal{Z},E,\mathcal{S})=-\delta\langle E,(\mathcal{O}-\mathcal{S})_{(3)}(\mathcal{Z}_{(3)})^{\top}\rangle$ of the form as (12b). Then the update of $E$ can be computed as follows
@@
where $\tilde{\alpha}_{E}=\frac{\delta}{\alpha_{E}}$ and $\operatorname{Proj}_{\mathbb{S}_{n_{3},r}}$ denotes the projection onto the Stiefel manifold $\mathbb{S}_{n_{3},r}$. It follows from Lemma 3.1 in [19] that the projection $\operatorname{Proj}_{\mathbb{S}_{n_{3},r}}$ has a closed form and $E^{k+1}$ can be computed as follows
@@ 3.3 The update of $\mathcal{Z}$
Then the subproblem for updating $\mathcal{Z}$ in (10) can be reformulated as
@@
${Z}^{k+1}_{::n}=\widetilde{\mathcal{D}}_{\sigma_{n}}^{\gamma}(\widehat{Z}^{k}_{::n}),$
@@ 4 Convergence Analysis of the PnP-PBCD Method
In this section, we conduct a convergence analysis on the proposed PnP-PBCD method. We first define the first-order optimality condition of problem (1) based on the subdifferentials for nonconvex nonsmooth functions and the Riemannian gradient of a smooth function on the Stiefel manifold. Then we prove that any accumulation point of the sequence generated by the PnP-PBCD method given in Algorithm 1 is a stationary point of problem (1).
4.1 First-order optimality condition
First, we provide some preliminaries on subdifferentials and Riemannian gradients.
@@ 4.1 First-order optimality condition
Second, we define the first-order optimality condition of the orthogonal constrained optimization problem as in (1). The point $(\bar{\mathcal{Z}},\bar{E},\bar{\mathcal{S}})$ is a first-order stationary point of problem (1) if
@@ 4.2 Subsequence convergence
    $\Phi_{\Sigma}$ is coercive and $\rho_{2}$-weakly convex.
@@
Let $f:\mathbb{R}^{d}\to(-\infty,+\infty]$ be a proper, lower semicontinuous and $\rho$-weakly convex function, and let $h:\mathbb{R}^{d}\to\mathbb{R}$ be a differentiable and $\mu$-strongly convex function. If there exists $\hat{x}$ such that
@@
where $\alpha\geq 0$, then
$f(x_{0})+h(x_{0})\geq f(\hat{x})+h(\hat{x})+\frac{\alpha+(\alpha+\mu-\rho)_{+}}{2}\|x_{0}-\hat{x}\|^{2},$ (20)
with $(\alpha+\mu-\rho)_{+}=\max\{\alpha+\mu-\rho,0\}$.
@@
For the $\rho$-weakly convex function $f$, it follows from Lemma 2.1 in [4] that
@@
for $\forall u\in\partial f(\hat{x})$. Similarly, for the differentiable and $\mu$-strongly convex function $h$, it follows from [21] that
@@
By the update of $\hat{x}$ in (19), we have $0\in\partial f(\hat{x})+\nabla h(\hat{x})+\alpha(\hat{x}-x_{0})$. That is, $-\nabla h(\hat{x})+\alpha(x_{0}-\hat{x})\in\partial f(\hat{x})$. Substituting $u=-\nabla h(\hat{x})+\alpha(x_{0}-\hat{x})$ into the above inequality, we obtain
$f(x_{0})+h(x_{0})\geq f(\hat{x})+h(\hat{x})+\frac{2\alpha+\mu-\rho}{2}\|x_{0}-\hat{x}\|^{2}.$
Note that by (19), we also have
$f(x_{0})+h(x_{0})\geq f(\hat{x})+h(\hat{x})+\frac{\alpha}{2}\|x_{0}-\hat{x}\|^{2}.$
Hence, taking the maximum of $2\alpha+\mu-\rho$ and $\alpha$, we obtain (20). ∎
Note that for $\alpha\geq 0$, it can be verified that $\alpha+(\alpha+\mu-\rho)_{+}>0$ if and only if $\alpha>0$ or $\mu-\rho>0$.
@@
Let $X\in\mathbb{R}^{n\times r}$, $A\in\mathbb{R}^{r\times m}$ and $B\in\mathbb{R}^{n\times m}$. Meanwhile, let $H_{1}(X)$, $H_{2}(X)$, and $H_{3}(X)$ denote the objective functions of (12a), (12b), and (12c), respectively. Then we have for any $X\in\mathbb{S}_{n,r}$,
@@
$\nabla H_{1}(X)$ $=(XA-B)A^{\top}=XAA^{\top}+\nabla H_{2}$
$\nabla H_{2}(X)$ $=-BA^{\top}$
$\nabla H_{3}(X)$ $=X-BA^{\top}=X+\nabla H_{2}.$
Then by projecting the gradients above onto the tangent space of Stiefel manifold at $X$, we can compute the Riemannian gradients of $H_{1}$, $H_{2}$, and $H_{3}$. Since $\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{n,r}}$ is linear, we have
$\operatorname{grad}H_{1}(X)$ $=\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{n,r}}(XAA^{\top})+\operatorname{grad}H_{2}(X)$
@@
$\operatorname{grad}H_{3}(X)$ $=\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{n,r}}(X)+\operatorname{grad}H_{2}(X).$
It is easy to verify that for any $X\in\mathbb{S}_{n,r}$,
$\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{n,r}}(XAA^{\top})=XAA^{\top}-\frac{1}{2}X(X^{\top}XAA^{\top}+(XAA^{\top})^{\top}X)=0$
and $\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{n,r}}(X)=X-\frac{1}{2}X(X^{\top}X+X^{\top}X)=0$. Therefore, we obtain (21). ∎
@@
Assume that assumptions (A1) and (A2) are satisfied. Let $F(\mathcal{Z},E,\mathcal{S})$ be the objective function of model (1) defined in (7) and let $\{(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\}$ be the sequence generated by Algorithm 1 with $\alpha_{\mathcal{S}}+(\alpha_{\mathcal{S}}+\delta-\rho_{1})_{+}>0$ and $\alpha_{\mathcal{Z}}+(\alpha_{\mathcal{Z}}+\delta-\rho_{2})_{+}>0$. Then the following statements hold:
@@
$F(\mathcal{Z}^{k-1},E^{k-1},\mathcal{S}^{k-1})-F(\mathcal{Z}^{k-1},E^{k-1},\mathcal{S}^{k})\geq\frac{\alpha_{\mathcal{S}}+(\alpha_{\mathcal{S}}+\delta-\rho_{1})_{+}}{2}\|\mathcal{S}^{k}-\mathcal{S}^{k-1}\|^{2}_{F}$
@@
$F(\mathcal{Z}^{k-1},E^{k},\mathcal{S}^{k})-F(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\geq\frac{\alpha_{\mathcal{Z}}+(\alpha_{\mathcal{Z}}+\delta-\rho_{2})_{+}}{2}\|\mathcal{Z}^{k}-\mathcal{Z}^{k-1}\|^{2}_{F}.$